T/F with explanations

1) The electric field inside a conducting wire causes electrons to drift left. This leads to a right flowing conventional current. Answer 2) Conventional current flows into a device at point A and out of the device at point B. The electric potential at point B is larger than the potential at point A Answer 3) The resistance of a wire is measured at an initial temperature. If the experiment was repeated at a higher temperature then the resistance of the wire would be smaller Answer 4) A battery is connected to an electronic device. If the connecting wires do not dissipate energy then the energy dissipated by the device equals the power of the battery. Answer 5) An AC generator causes the current flowing in an electronic device to periodically reverse direction. The peak value of the current is larger than the rms value of the current. Answer
Answer #1

Solution) (1) True

Usually in conductors (metals) charge carriers (electrons) are negative and conventional current is in opposite direction as electrons

(2) False

Conventional current flows from high potential to low potential

(3) False

Rt = Ro(1 + a(dT))

Here Rt is resistance at higher temperature

Ro is initial temperature

a(alpha) is coefficient of expansion

dT is change in temperature

So at higher temperature resistance increases

(4) TRUE

Battery Power P = Vi = (iR)i = (i^2)R = power dissipated

(5) TRUE

We have

Ip = Irms × (2)^(1/2)

Ip is peak current

Irms is rms current

So from above relation peak current is larger than rms current